# polyGen - A Learning Framework for Atomic-level Polymer Structure Generation

### v1

@@ Abstract
Synthetic polymeric materials underpin fundamental technologies in the energy, electronics, consumer goods, and medical sectors, yet their development still suffers from prolonged design timelines. Although polymer informatics tools have supported speedup, polymer simulation protocols continue to face significant challenges: on-demand generation of realistic 3D atomic structures that respect the conformational diversity of polymer structures. Generative algorithms for 3D structures of inorganic crystals, bio-polymers, and small molecules exist, but have not addressed synthetic polymers. In this work, we introduce polyGen, the first latent diffusion model designed specifically to generate realistic polymer structures from minimal inputs such as the repeat unit chemistry alone, leveraging a molecular encoding that captures polymer connectivity throughout the architecture. Due to a scarce dataset of only 3855 DFT optimized polymer structures, we augment our training with DFT-optimized molecular structures, showing improvement in joint learning between similar chemical structures. We also establish structure matching criteria to benchmark our approach on this novel problem. polyGen effectively generates diverse conformations of both linear chains and complex branched structures, though its performance decreases when handling repeat units with a high atom count. Given these initial results, polyGen represents a paradigm shift in atomic level structure generation for polymer science—the first proof-of-concept for predicting realistic atomic-level polymer conformations while accounting for their intrinsic structural flexibility.
@@ 1 Introduction
The complexity of polymer design and representation, combined with limited available data, explains why generative models for polymers have remained largely unexplored until now. We introduce polyGen (Figure 2), the first latent diffusion model specifically designed to generate periodic and stochastic polymeric structures. We build upon latent diffusion frameworks for materials generation such as All-Atom DiT [30]. Our problem and approach are modified to predict an ensemble of low-energy polymer conformations conditioned on a specified repeat unit with prescribed atomic connectivities rather than performing unconstrained de novo structure generation. Our method leverages a molecular encoding that captures atomic connectivity, which is used as a conditioning signal throughout the model architecture. We adopt a latent diffusion strategy over joint diffusion methods such as DiffCSP [18] or molecular generation methods in Cartesian space [23] due to the strong coupling between atomic positions and box dimensions imposed by polymer connectivity constraints. We augment our dataset with smaller structures from DFT-optimized molecules, demonstrating improved polymer structure prediction capabilities owing to the shared weights of a dataset invariant encoder and decoder, and a shared latent space. This work culminates in a learning framework that can generate ensembles of realistic polymer chain conformations. To benchmark the quality of these conformations, we introduce rigorous evaluation criteria grounded in first principles—standards that, to our knowledge, have not yet been applied in the context of materials generation. This application represents a novel proof-of-concept to predict atomic-level synthetic polymer conformations while accounting for their intrinsic flexibility and the distribution of plausible structures.
@@ 2.1 Dataset and Evaluation Metrics
We benchmark polymer structure prediction on our polyChainStructures test set with our method. Unlike crystals, polymers are amorphous and can adopt many valid low-energy conformations, making conventional structure matching [32] unsuitable. Molecule generation is also evaluated with root mean squared distance after rotational alignment, but this doesn’t account for conformational variability experienced by infinite chains.
@@ 2.2 Structure Matching Results
We find that our model is capable of predicting polymer structures that closely match the linear chain structure calculated by DFT. To provide context to our rationale, Figure 3 shows examples of predicted polymers compared to the ground truth of the dataset. In Figure 3a-d we see the capabilities of generating systems with qualitatively similar structures after filtering. We also provide examples of errors that are caught by our filtering. In Figure 3a and c, we see examples of carbon atoms incorrectly predicted within a supposedly aromatic ring in the backbone, which lead to incorrect connectivity. Despite Figure 3b being a failed structure, this error can easily be fixed with an energy minimization or a heuristic-based increasing of the C-H bond, demonstrating the efficacy of the remaining generation and the stringent nature of the generation filter. Also, the complexity of these structures should be noted, given chain size, branching, and the number of rings present.
@@
Figure 3: Visual samples from 100 generations per polymer type. Each row displays examples from a single polymer, accompanied by corresponding ground truth distributions (top) and model-predicted distributions (bottom) of molecular bonds, angles, and dihedrals for that polymer. For a-c, the ground truth, a successfully generated sample, and an unsuccessful sample are provided. The unsuccessful samples highlight the errors that caused the incorrect connectivity. Note that the structures for b) and c) are replicated along the z-axis for visual clarity of the structure. d) An example of a simple repeat unit for which no predictions fail. e) An example of the largest structure in the test set, with 208 atoms, for which all generations fail.
@@
To quantify structural similarity, we compare bond length, triplet angle, and dihedral angle distributions for specific atomic species. Like X-ray diffraction analysis, predicted structures should exhibit peaks comparable to optimized structures. On the right side of Figure 3, we observe variable distributions—most predictions cluster near optimized peaks with some outliers. The forward KL divergences establish benchmarks for acceptable "matches" in bond lengths, angles, and dihedrals.
@@
We find that our approach can find relative trends in the majority of bond lengths, but lacks precision. For example, the predicted C-O bonds in Figure 3a and C-C bonds in Figure 3d show 2 peaks for single and double bonds, with noisy predictions scattered by at least 0.1 Å around these peaks. Additionally, the distribution of the predicted C-S bonds in Figure 3c includes an appropriate higher bond length, but with a much higher dispersity of predictions which results in a high KL divergence of 3.08. A similar example is found for the C-N single bond in Figure 3b. This makes it evident that generation is precise to the order of Å, but not on the scale of picometers. Precision on the level of picometers will be necessary to properly distinguish between bond types before our approach is scaled to larger system sizes.
Angles and dihedrals might vary based on the conformation. We find that the nature of these distributions also matches the ground truth peaks, especially the O-C-O, O-S-O, and C-C-C bonds in Figure 3a, c and d respectively. The dihedral distributions can be used to validate the prediction of correct local structures in the polymer. For example, the C-N-C-C dihedrals in Figure 3b validate the feasible prediction of the nitrogen containing 5-membered ring in the backbone. In Figure 3c, the C-C-C-O dihedral on the long branch is predicted similarly in a majority of cases based on the probability density, but has a wider extraneous variability of predictions. This conveys the challenges in predicting structures for branched polymers.
Figure 3e shows the failed predictions for the largest system across the train, validation, and test datasets, which is in our test set. Since none of the generations pass our filter, we show distributions from all failed predictions. We see the wide distribution of C-C bonds which extends to <1.0 Å due to the irregular placement of C atoms in these failed structures. A similar occurrence can be seen in the predicted C-C-N angle, which matches the peak of the DFT optimized structure, but with a larger distribution of noisy peaks. These results place a clear limit on the size of system due to the training dataset constraints, demonstrating the challenges of predicting feasible structures for larger polymer repeat units.
@@ 2.4 Overall Results
As highlighted in Figure 6d and discussed in Section 2.3, we expect some variability in the prediction of the z-height of the polymer chain. Despite this, we can achieve an $r^{2}$ score of 0.822 between the predicted and DFT heights, meaning that the model can differentiate between dense vs. sparsely packed structures. Additionally, we see a larger standard deviation and error in predictions for larger repeat units (> 10 Å), because these repeat units may show more flexibility. Figure 6E also shows a notable trend in the difficulty of generating feasible structures as system size increases. Most generated systems with <25 atoms show success rates > 0.5. For system sizes greater than 150 atoms, we find errors in connectivity for all generated systems, similar to the examples shown in Figure 3E. Intuitively, it would also seem that complex polymeric structures would be more difficult to generate. To this end, we also compare these metrics with the Synthetic Accessibility (SA) in Supplementary Information A. We see loose correlations with KL divergences of bond length, angles, and dihedrals with the SA score, but the ratio of successful generations of a polymer do not have a strong correlation with the SA score. Therefore, generation success could be independent from SA score and more reliant on the system size.
Figure 5: KL Divergences of Joint training vs. polyChainStructures only for a) Bonds b) Angles and c) Dihedrals. d) The comparison of z-height prediction vs the DFT ground truth. e) The success rate of filtering as a function of the atom count of the system. A histogram in the back provides the distributions of atom counts found in the test dataset, with the density provided on the right y-axis.
Methodologically, increasing the diffusion transformer’s size [30] and addressing the sign-invariance of Laplacian eigenvector positional encodings—using techniques such as SignNet [33, 34]—could further enhance the framework’s performance especially with larger systems.
@@ 3 Discussion
In this study, we introduce a solution to the problem of atomic-level polymer structure generation—given only the atomic connectivity of a repeat unit (i.e. SMILES), polyGen can generate an ensemble of realistic 3D structures of synthetic polymers. Results demonstrate that polyGen effectively generates structures with bond lengths, angles, and dihedral distributions that align well with ground truth structures. The model also successfully generates valid conformations for complex features like aromatic backbones and branches while correlating repeat unit chemistry with structural properties, such as repeat unit length. These initial results are particularly promising considering our polymer structure dataset optimized with Density Functional Theory contains only 3855 systems. We also investigate the limitations and demonstrate the need for stringent benchmarking of future polymer structure generation techniques. While capturing relative trends in structure, the model lacks precision at the picometer scale needed to definitively distinguish between bond types. Performance degrades significantly for larger polymeric systems, likely due to dataset constraints. Finally, the consistency of feasible generation needs to be improved. Given the successes of this proof-of-concept, future work will focus on expanding the training dataset to include larger polymer systems, incorporating additional physics-informed constraints, and exploring hybrid approaches that combine latent diffusion with molecular dynamics simulations. Addressing these challenges could transform polyGen into an invaluable tool for computational materials science, accelerating the discovery of novel polymeric materials across numerous applications.
@@ 4.2 Graph Conditioning
[…]
$\mathbf{h}_{i}^{(0)}$ $=\text{MLP}([\mathbf{h}_{i}^{(0)};\mathbf{r}_{i}^{\text{RW}};\mathbf{r}_{i}^{\text{Lap}}])$
For each layer $l=1,\ldots,L$:
Edge Update:
@@
Node Update:
$\mathbf{m}_{i}^{(l)}$ $=\sum_{j\in\mathcal{N}(i)}\mathbf{e}_{ij}^{(l)}$ (4)
$\mathbf{h}_{i}^{(l)}$ $=\mathbf{h}_{i}^{(l-1)}\mathbf{W}_{v}+f_{v}^{(l)}(\text{LN}([\mathbf{m}_{i}^{(l)};\mathbf{h}_{i}^{(l-1)}]))$
Global Pooling and Final Representation:
$\mathbf{g}$ $=\sum_{i\in\mathcal{V}}\mathbf{h}_{i}^{(L)}$ (5)
$C_{i}$ $=f_{f}^{(l)}([\mathbf{g};\mathbf{h}_{i}^{(L)}])$
@@ 4.3 Structural Variational AutoEncoder
$\mathcal{L}_{\text{bbox}}$ $=\text{MSE}({\hat{b}}_{z},b_{z}/(10*\sqrt[3]{N}))$
$\mathcal{L}_{\text{frac\_coords}}$ $=\frac{1}{d}\sum_{i}\text{MSE}(\hat{f}_{i},f_{i})$
$\mathcal{L}_{\text{pos}}$ $=\frac{1}{d}\sum_{i}\text{MSE}((\hat{p}_{i}-\langle\hat{p}\rangle),(p_{i}-\langle p\rangle))$
$\mathcal{L}_{\text{kl}}$ $=\text{KL}(q_{\phi}(z|x)\|p(z))$
And the structural loss terms from are:
$\mathcal{L}_{\text{bond}}$ $=\text{MSE}(d_{ij},\hat{d}_{ij})$
$\mathcal{L}_{\text{angle}}$ $=\text{MSE}(\angle_{ijk},\hat{\angle}_{ijk})$
$\mathcal{L}_{\text{dihedral}}$ $=(\Delta_{\text{periodic}}(\tau_{ijkl},\hat{\tau}_{ijkl}))^{2}$
@@
$d_{ij}$ $=\text{bond length between atoms $i$ and $j$ computed with true coordinates and box}$
$\hat{d}_{ij}$ $=\text{bond length between atoms $i$ and $j$ computed with predicted coordinates and box}$
$\angle_{ijk}$ $=\text{angle between bonded atoms $i$-$j$-$k$ computed with true coordinates}$
^$\hat{\angle}_{ijk}$ $=\text{angle between bonded atoms $i$-$j$-$k$ computed with predicted coordinates}$
$\tau_{ijkl}$ $=\text{dihedral angle formed by atoms $i$-$j$-$k$-$l$ computed with true coordinates}$
$\hat{\tau}_{ijkl}$ $=\text{dihedral angle formed by atoms $i$-$j$-$k$-$l$ computed with predicted coordinates}$
@@ 4.4 Diffusion Transformer
Now that a joint latent space is learned, we require a way to find suitable structures for new polymer chemistries. We use a DiT architecture for our generative model $\mathcal{M}$, operating within the latent space $\mathcal{Z}$ learned by the VAE. We use a similar DiT architecture as ADiT [30], however the previous work uses sinusoidal positional encodings for atomic tokens, which make it difficult for unordered atomic representations. In our case, we want to preserve the permutation invariant qualities of a GNN, so we use the positional encoded features and interactions in $C_{i}$ to provide positional information relative to other atoms in the system.
@@
The final prediction task for the graph diffusion model is defined as,
$\hat{\mathcal{Z}}_{1}$ $=\mathcal{M}(\mathcal{Z}_{t},\mathcal{Z}_{\text{sc}},t,C_{i},S)$
$\hat{u}_{t}$ $=\frac{\hat{\mathcal{Z}}_{1}-\mathcal{Z}_{t}}{1-t},$
where $S$ is an embedding that represents the dataset used during generation. This is used exclusively during joint training. Only the embedding for the polyChainStructures dataset is used during inference. $\mathcal{Z}_{\text{sc}}$ denotes a self-conditioning input, which corresponds to a previous prediction of $\hat{\mathcal{Z}}_{1}$. Self-conditioning has been shown to improve performance in autoregressive molecular generation tasks [41, 30].
In our training implementation, we follow a two-pass approach:
  • In the first pass, we predict $\hat{\mathcal{Z}}_{1}$ using no self-conditioning, i.e., $\mathcal{Z}_{\text{sc}}=\emptyset$.
  • In the second pass, the predicted $\hat{\mathcal{Z}}_{1}$ is fed back as $\mathcal{Z}_{\text{sc}}$ to provide additional context for refinement.
To regularize the model and prevent over-reliance on self-conditioning, we follow standard practice and randomly drop the self-conditioning input during training [41, 30]. Specifically, with probability 0.5, we set $\mathcal{Z}_{\text{sc}}=\emptyset$.

### v2

@@ Abstract
Synthetic polymeric materials underpin fundamental technologies in the energy, electronics, consumer goods, and medical sectors, yet their development still suffers from prolonged design timelines. Although polymer informatics tools have supported speedup, polymer simulation protocols continue to face significant challenges in the on-demand generation of realistic 3D atomic structures that respect the conformational diversity of polymers. Generative algorithms for 3D structures of inorganic crystals, bio-polymers, and small molecules exist, but have not addressed synthetic polymers because of challenges in representation and dataset constraints. In this work, we introduce polyGen, the first generative model designed specifically for polymer structures from minimal inputs such as the repeat unit chemistry alone. polyGen combines graph-based encodings with a latent diffusion transformer using positional biased attention for realistic conformation generation. Given the limited dataset of 3,855 DFT-optimized polymer structures, we incorporate joint training with small molecule data to enhance generation quality. We also establish structure matching criteria to benchmark our approach on this novel problem. polyGen overcomes the limitations of traditional crystal structure prediction methods for polymers, successfully generating realistic and diverse linear and branched conformations, with promising performance even on challenging large repeat units. As the first atomic-level proof-of-concept capturing intrinsic polymer flexibility, it marks a new capability in material structure generation.
@@ 1 Introduction
The complexity of polymer design and representation, combined with limited available data, explains why generative models for polymers have remained largely unexplored until now. We introduce polyGen (Figure 2), the first latent diffusion model specifically designed to generate periodic and stochastic polymeric structures. We build upon latent diffusion frameworks for materials generation such as All-Atom DiT [30]. Our problem and approach are modified to predict an ensemble of low-energy polymer conformations conditioned on a specified repeat unit with prescribed atomic connectivities rather than performing unconstrained de novo structure generation. Our method leverages a molecular encoding that captures atomic connectivity, which is used as a conditioning signal throughout the model architecture. We adopt a latent diffusion strategy over joint diffusion methods such as DiffCSP [18] or molecular generation methods in Cartesian space [23] due to the strong coupling between atomic positions and box dimensions imposed by polymer connectivity constraints. We augment our dataset with smaller structures from DFT-optimized molecules, demonstrating improved polymer structure prediction capabilities owing to the shared weights of a dataset invariant encoder and decoder, and a shared latent space. This work culminates in a learning framework that can generate ensembles of realistic polymer chain conformations. To benchmark the quality of these conformations, we introduce rigorous evaluation criteria on bonds, angles, and dihedrals—standards that, to our knowledge, have not yet been applied in the context of materials generation. This application represents a novel proof-of-concept to predict atomic-level synthetic polymer conformations while accounting for their intrinsic flexibility and the distribution of plausible structures.
@@ 2.1 Dataset and Evaluation Metrics
We benchmark polymer structure prediction on our polyChainStructures test set with our method. Unlike crystals, polymers are amorphous and can adopt many valid low-energy conformations, making conventional structure matching [32] unsuitable. Molecule generation is also evaluated with root mean squared distance after rotational alignment [24, 33], but this doesn’t account for conformational variability experienced by infinite chains.
@@ 2.2 Structure Matching Results
We find that our model is capable of predicting polymer structures that closely match the linear chain structure calculated by DFT. To provide context to our rationale, Figure 3 shows examples of predicted polymers compared to the ground truth of the dataset. In Figure 3a-d we see the capabilities of generating systems with qualitatively similar structures after filtering. We also provide examples of errors that are caught by our filtering. In Figure 3a and c, we see examples of carbon atoms incorrectly predicted within a supposedly aromatic ring in the backbone, which lead to incorrect connectivity. Many failed structures, especially the one in Figure 3b, can easily be fixed with an energy minimization or a heuristic-based increasing of the C-H bond, demonstrating the efficacy of the remaining generation and the stringent nature of the generation filter. Also, the complexity of these structures should be noted, given chain size, branching, and the number of rings present.
@@
Figure 3: Visual samples from 100 generations per polymer type. Each row (a-d) displays examples from a single polymer. The ground truth, a successfully generated sample, and an unsuccessful sample are provided. The unsuccessful samples highlight the errors that caused the incorrect connectivity. Note that the structures for b) and c) are replicated along the z-axis for visual clarity of the structure. The visualizations are accompanied by corresponding ground truth distributions (top) and model-predicted distributions (bottom) of molecular bonds, angles, and dihedrals for that polymer.
@@
To quantify structural similarity, we compare bond length, triplet angle, and dihedral angle distributions for some atomic species. Predicted structures should exhibit peaks comparable to optimized structures. The forward KL divergences establish benchmarks for acceptable "matches" in bond lengths, angles, and dihedrals.
@@
We find that our approach can find relative trends in the majority of bond lengths, but lacks precision. For example, the predicted C-O bonds in Figure 3a shows 2 peaks for single and double bonds, with noisy predictions scattered by at least 0.1 Å around these peaks. In Figure 3b and c, both predicted C-N and C-S bond distributions show peaks corresponding with the true bond length, but with wide distributions. Generation is precise to the order of Å, but not on the scale of picometers. Precision on the level of picometers will be necessary to properly distinguish between bond types before our approach is scaled to larger system sizes.
[…]
Figure 3d shows predictions for the largest system in the test dataset, containing 208 atoms. Only 3 out of 100 generated structures pass the filtering when using position-biased attention, while the DiT model with vanilla attention fails to produce any valid structure. The broad distribution of predicted C–N–C–O dihedral angles and their high deviation from the ground truth further highlight the lack of conformational viability. These results underscore a clear limitation in handling larger systems, likely due to constraints in the training dataset, and illustrate the challenges of generating feasible structures for complex polymer repeat units.
@@ 2.4 Overall Results
Figure 5: KL Divergences of Joint training for Position Bias, Joint Training and polyChainStructures Only for a) Bonds b) Angles and c) Dihedrals. d) The comparison of z-height prediction vs the DFT ground truth. e) The success rate of filtering as a function of the atom count of the system.
@@
Another major improvement is seen when including position attention biasing (Section 4.4.1) in the DiT module. Because local interactions are explicitly specified, we find that the positional encoding improves generation success from 36.9% to 64.8% when looking at isomorphism of the predicted structure compared to the original graph. Figure 5 showcases the improvement of polyGen with a relative encoding bias on all metrics, notably the decrease of angle and dihedral KL Divergences by 33.5% and 21.2%, respectively.
As highlighted in Figure 5d and discussed in Section 2.3, we expect some variability in the prediction of the z-height of the polymer chain. Despite this, we can achieve an $r^{2}$ score of 0.854 between the predicted and DFT heights, meaning that the model can differentiate between dense vs. sparsely packed structures. Additionally, we see a larger standard deviation and error in predictions for larger repeat units (> 10 Å), because these repeat units may show more flexibility. Figure 5e also highlights the difficulty of generation as system size increases. Most generated systems with <25 atoms show success rates > 0.5. For system sizes greater than 150 atoms, the model with vanilla attention cannot generate any systems with proper connectivity, but the relative position bias substantially improves this.
Intuitively, it would also seem that complex polymeric structures would be more difficult to generate. To this end, we also compare these metrics with the Synthetic Accessibility (SA) in Supplementary Information A. We see loose correlations with KL divergences of bond length, angles, and dihedrals with the SA score, but the ratio of successful generations of a polymer do not have a strong correlation with the SA score. Therefore, generation success could be independent from SA score and more reliant on the system size.
@@ 3 Discussion
In this study, we introduce a solution to the problem of atomic-level polymer structure generation—given only the atomic connectivity of a repeat unit (i.e. SMILES), polyGen can generate an ensemble of realistic 3D structures of synthetic polymers. Results demonstrate that polyGen effectively generates structures with bond lengths, angles, and dihedral distributions that align well with ground truth structures, and the quality of structures improves with the inclusion of positional biasing to the attention mechanism of the diffusion transformer. The model also successfully generates valid conformations for complex features like aromatic backbones and branches while correlating repeat unit chemistry with structural properties, such as repeat unit length. These initial results are particularly promising considering our polymer structure dataset optimized with Density Functional Theory contains only 3855 systems. We also investigate the limitations and demonstrate the need for stringent benchmarking of future polymer structure generation techniques. While capturing relative trends in structure, the model lacks precision at the picometer scale needed to definitively distinguish between bond types. Performance degrades significantly for larger polymeric systems, likely due to dataset constraints. Finally, the consistency of feasible generation needs to be improved. Given the successes of this proof-of-concept, future work will focus on expanding the training dataset to include larger polymer systems, incorporating additional physics-informed constraints, and exploring hybrid approaches that combine latent diffusion with molecular dynamics simulations. Addressing these challenges could transform polyGen into an invaluable tool for computational materials science, accelerating the discovery of novel polymeric materials across numerous applications.
@@ 4.2 Graph Conditioning
For each layer $l=1,\ldots,L$, we update the edges and the nodes as:
@@
$\mathbf{m}_{i}^{(l)}$ $=\sum_{j\in\mathcal{N}(i)}\mathbf{e}_{ij}^{(l)},\quad\mathbf{h}_{i}^{(l)}=\mathbf{h}_{i}^{(l-1)}\mathbf{W}_{v}+f_{v}^{(l)}(\text{LN}([\mathbf{m}_{i}^{(l)};\mathbf{h}_{i}^{(l-1)}]))$
We finally node features to add global context for the whole molecule:
$\mathbf{g}$ $=\sum_{i\in\mathcal{V}}\mathbf{h}_{i}^{(L)},\quad C_{i}=f_{f}^{(l)}([\mathbf{g};\mathbf{h}_{i}^{(L)}])$ (4)
@@ 4.3 Structural Variational AutoEncoder
$\mathcal{L}_{\mathrm{bbox}}$ $=\mathrm{MSE}\bigl{(}\hat{b}_{z},\;b_{z}/(10\sqrt[3]{N})\bigr{)},$ $\mathcal{L}_{\mathrm{frac\_coords}}$ $=\frac{1}{d}\sum_{i}\mathrm{MSE}(\hat{f}_{i},f_{i}),$
$\mathcal{L}_{\mathrm{pos}}$ $=\frac{1}{d}\sum_{i}\mathrm{MSE}\bigl{(}\hat{p}_{i}-\langle\hat{p}\rangle,\;p_{i}-\langle p\rangle\bigr{)},$ $\mathcal{L}_{\mathrm{kl}}$ $=\mathrm{KL}\bigl{(}q_{\phi}(z\mid x)\;\|\;p(z)\bigr{)}.$
And the structural loss terms are:
$\mathcal{L}_{\text{bond}}$ $=\text{MSE}(d_{ij},\hat{d}_{ij})$ $\mathcal{L}_{\text{angle}}$ $=\text{MSE}(\angle_{ijk},\hat{\angle}_{ijk})$ $\mathcal{L}_{\text{dihedral}}$ $=(\Delta_{\text{periodic}}(\tau_{ijkl},\hat{\tau}_{ijkl}))^{2}$
@@
[…]
@@ 4.4 Diffusion Transformer
Now that a joint latent space is learned, we require a way to find suitable structures for new polymer chemistries. We use a DiT architecture for our generative model $\mathcal{M}$, operating within the latent space $\mathcal{Z}$ learned by the VAE. We use a similar DiT architecture as ADiT [30], however the previous work uses sinusoidal positional encodings for atomic tokens, which make it difficult for unordered atomic representations. In our case, we want to preserve the permutation invariant qualities of a GNN, so we use the positional encoded features and interactions in $C_{i}$ to provide positional information relative to other atoms in the system. We also modify the attention mechanism to add a bias towards bonds, elaborated in Section 4.4.1.
@@
The final prediction task is defined as,
$\hat{\mathcal{Z}}_{1}=\mathcal{M}(\mathcal{Z}_{t},\mathcal{Z}_{\text{sc}},t,C_{i},S),\quad\hat{u}_{t}=\frac{\hat{\mathcal{Z}}_{1}-\mathcal{Z}_{t}}{1-t}$
where $S$ is an embedding that represents the dataset used during generation. This is used exclusively during joint training. Only the embedding for the polyChainStructures dataset is used during inference. $\mathcal{Z}_{\text{sc}}$ denotes a self-conditioning input, which corresponds to a previous prediction of $\hat{\mathcal{Z}}_{1}$. Self-conditioning improves autoregressive molecular generation [41, 30]. Our training uses a two-pass approach: first predicting $\hat{\mathcal{Z}}_{1}$ with $\mathcal{Z}_{\text{sc}}=\emptyset$, then feeding this prediction back as $\mathcal{Z}_{\text{sc}}$ for refinement. To avoid overreliance, we randomly drop self-conditioning ($\mathcal{Z}_{\text{sc}}=\emptyset$) with probability 0.5 during training. Within $\mathcal{M}$, the latent features and conditioning signals are incorporated as
$\tilde{x}=E_{\mathcal{Z}}([\mathcal{Z}_{t};\mathcal{Z}_{\text{sc}}])+C_{i},\quad c=d+E_{t}(t),$
where $\tilde{x}$ is input and $c$ is the modulation conditioning for attention blocks from [Peebles2022DiT]. $E_{\mathcal{Z}}$ and $E_{t}$ are embedding blocks.
@@
4.4.1 Relative Positional Encoding Attention Modification
With a traditional attention mechanism, DiT is forced to learn bonding relationships from the atom-wise conditional embeddings. For smaller systems, this can provide enough information for proper structure prediction, but this may be a problem for larger systems. Alphafold3, which operates on larger atomic systems than traditional material structure generation models, utilizes a pairformer [21] with a relative position encoding to allow tokens to attend to nearby neighbors by conditioning the attention weights of the diffusion module.
We employ a lightweight relative positional encoding with attention biasing mechanism in the polyGen DiT module to differentiate local atomic interactions from global. First, we encode pairwise relationships with a one-hot graph-distance tensor
$D\;\in\;\{0,1\}^{N\times N\times 5},$
where each of the five channels indicates whether atoms $i$ and $j$ are
  1. (1)
    identical
  2. (2)
    bonded
  3. (3)
    separated by an angle
  4. (4)
    separated by a dihedral
  5. (5)
    beyond four bonds apart
An MLP $f_{D}$ then maps each one-hot vector $D_{i,j}$ to a scalar bias, which is added directly into the scaled dot-product attention:
$\mathrm{bias}_{i,j}$ $=f_{D}\bigl{(}D_{i,j}\bigr{)},$ (14)
$\mathrm{Attention}_{i,j}$ $=\mathrm{Softmax}\!\Bigl{(}\frac{Q_{i}\cdot K_{j}}{\sqrt{\mathrm{num\_heads}}}+\mathrm{bias}_{i,j}\Bigr{)}.$ (15)
Intuitively, nearby atoms (bonds, angles, dihedrals) should have a higher bias and other atoms should have a lower bias. With a per-DiT-block learnable bias, the DiT has the ability to allocate more attention weight while still retaining the ability to attend globally in some layers.